A test consists of 10 multiple choice questions with five choices for each question. As an experiment,
you GUESS on each and every answer without even reading the questions.
What is the probability of getting exactly 7 questions correct on this test?

Answer by stanbon(75887)

A test consists of 10 multiple choice questions with five choices for each question. As an experiment,
you GUESS on each and every answer without even reading the questions.
What is the probability of getting exactly 7 questions correct on this test?
------
Binomial Problem with n = 10 and p(correct) = 1/5
-------------------
P(x = 7) = 10C7*(1/5)^7*(4/5)^3 = binompdf(10,1/5,7) = 0.000786..
-----
Cheers,
Stan H.
--------------------